OpenStudy (anonymous):

What is the probability of rolling a 6 twice in a row with a number cube that has the numbers 1, 2, 3, 4, 5, and 6? A. B. C. D.

OpenStudy (anonymous):

the probability of getting a 6 on the first roll is 1/6, same with the second roll. so to get it rolling both times is 1/6 * 1/6 = 1/36.

OpenStudy (perl):

P( A & B ) = P(A) * P(B) , assuming A and B are independent

OpenStudy (anonymous):

oh ok... thank you!

OpenStudy (anonymous):

yw